# Multilinear Wavelet Compact T(1) Theorem

### v1

@@ A CZ form $\Lambda$ is said to be a compact CZ form if for some (all) $M\geq 1$
It is not too difficult to see that if $\mathsf{W}_{\Lambda}^{M}\to 0$ for some $M$ then the same holds for each $M$. Furthermore, since ${\mathcal{S}}^{0}(\mathbb{R}^{d})$ is dense in the Hardy space $H^{1}(\mathbb{R}^{d})$, which is the dual space of $\mathrm{BMO}(\mathbb{R}^{d})$, (3.1) should be interpreted as a weak limit, i.e. $T^{*,j}(\theta_{t},\ldots,\theta_{t})$ converges weakly to $\mathfrak{b}_{j}$ in $\mathrm{BMO}(\mathbb{R}^{d})$.
@@
Let $T$ be an $m$-linear CZO. We say $T$ is a compact CZO if for each $(\vec{p},r)\in Q$, $T(\mathcal{B}^{\vec{p}})$ is precompact in $L^{r}(\mathbb{R}^{d})$, and at the upper endpoint, $T(\mathcal{B}^{\infty}_{0})$ is precompact in $\mathrm{CMO}(\mathbb{R}^{d})$, where
$\mathcal{B}^{\infty}_{0}=\left\{(f_{1},\ldots,f_{m})\in\mathcal{B}^{(\infty,\ldots,\infty)}:f_{j}\ \mathrm{compactly}\ \mathrm{supported}\right\}.$
@@
Let $\Lambda$ be an $(m+1)$-linear CZ form such that each $T\in\mathsf{T}_{\Lambda}$ is a compact CZO. Then, each $\mathfrak{b}_{j}\in\mathrm{CMO}(\mathbb{R}^{d})$, $j=0,\ldots,m$, and for each $M\geq 1$,
@@
To prove the first conclusion, since $\Lambda$ is a CZ form, (3.1) holds. Let us fix $j=0,\ldots,m$. Up to an absolute constant, $(\theta_{t},\ldots,\theta_{t})\in\mathcal{B}^{\infty}_{0}$, so there exists $t_{n}\to 0$ such that $T^{*,j}(\theta_{t_{n}},\ldots,\theta_{t_{n}})$ converges in $\mathrm{CMO}(\mathbb{R}^{d})$. However, the second statement in Remark 3.3 requires this limit to coincide with $\mathfrak{b}_{j}$, and the first part is proved. To prove the second conclusion, let us suppose, towards a contradiction, that there exists $M\geq 1$ such that (3.4) does not hold. Therefore we can find $\sigma\in\mathrm{S}_{m+1}$, $\vec{\psi}\in{\mathcal{S}}(\mathbb{R}^{d})^{m}$, $\varepsilon>0$ and sequences $\zeta_{n}=(w_{n},t_{n})\in Z^{d}$ and $\mathbf{z}_{n}\in Z^{md}$ such that
@@
(3.5) $\zeta_{n}\to\infty,\quad\mathbf{z}_{n}\in\mathbb{B}^{m}_{M}(\zeta_{n}),\quad\left|\Lambda^{\sigma}\left(\phi_{n},\vec{\psi}_{n}\right)\right|\geq\varepsilon,\quad\phi_{n}=t_{n}^{\frac{d}{r}}\phi_{\zeta_{n}},\quad\vec{\psi}_{n}=t_{n}^{d(m-\frac{1}{r})}\vec{\psi}_{\mathbf{z}_{n}},$
and $r>1$. Let us suppose, for simplicity that $\sigma$ is the identity. If not, then in what follows we would replace $T\vec{\psi}_{n}$ by $T^{*,\sigma(0)}(\vec{\psi}^{\sigma^{\prime}}_{n})$ where $\vec{\psi}^{\sigma^{\prime}}_{n}$ is an appropriate permutation of the $m$ elements of $\vec{\psi}_{n}$. Since $\{\vec{\psi}_{n}\}_{n\in\mathbb{N}}$ is a bounded sequence in $L^{\vec{p}}(\mathbb{R}^{d})$, for some $\vec{p}\in P$ with $r(\vec{p})=r$, by the compactness of $T$ we find a subsequence so that $\{T\vec{\psi}_{n_{k}}\}_{k\in\mathbb{N}}$ is convergent in $L^{r}(\mathbb{R}^{d})$. Finally, it is not hard to check that since $\zeta_{n}\to\infty$, $\phi_{n}$ converges weakly to zero in $L^{\frac{r}{r-1}}(\mathbb{R}^{d})$ which, since $r>1$, is the dual space of $L^{r}(\mathbb{R}^{d})$. Therefore,
$\Lambda^{\sigma}(\phi_{n_{k}},\vec{\psi}_{n_{k}})=\left\langle\phi_{n_{k}},T\vec{\psi}_{n_{k}}\right\rangle\to 0,\,\mbox{as}\,k\to\infty,$
which contradicts (3.5). ∎
@@
$\left|\Lambda\left(\vec{\psi}_{\mathbf{z}},\phi_{\zeta}\right)\right|\lesssim\left(t^{-md}\right)^{1-\theta}\left[\mathbf{z},\zeta\right]_{\delta_{2}}^{\theta}\mathsf{W}_{\Lambda}^{n}(\zeta)^{1-\theta},\quad\mathbf{z}\in\mathbb{B}^{m}_{n}(\zeta),\ \mathbf{z}\geq\zeta.$
We choose $\theta$ close enough to $1$ that $md\theta+\delta_{2}\theta-md=\delta_{1}$. With this specific choice of $\theta$, one can easily verify that $\left(t^{-md}\right)^{1-\theta}[\mathbf{z},\zeta]_{\delta_{2}}^{\theta}=[\mathbf{z},\zeta]_{\delta_{1}}$. Setting $\varepsilon_{z}^{1}=\mathsf{W}^{1}_{\Lambda}(\zeta)^{1-\theta}$ handles the case $n=1$. For $n\geq 2$, we factor, with $\rho=\delta_{1}-\eta$,
@@
Therefore, one only needs to verify that for $\mathbf{z}\in\mathbb{B}^{m}_{n}(\zeta)\setminus\mathbb{B}^{m}_{{n-1}}(\zeta)$, the first factor is comparable to $2^{-\rho n}$. The proof is concluded by setting $\varepsilon^{n}_{\zeta}=\mathsf{W}^{n}_{\Lambda}(\zeta)$. ∎
@@
We recycle the proof of the representation theorem in [DGW] relying on Lemma 2.4 and Lemma 2.5. First expand each $f_{j}$ using (2.3) to obtain
@@ 4. Compact wavelet forms
Let $U$ be a compact wavelet form. Then each $T\in\mathsf{T}_{U}$ is a compact CZO.
@@
Let us fix a compact wavelet form $U$, $T\in\mathsf{T}_{U}$, and $(\vec{p},r)\in Q$. $T(f_{1},\ldots,f_{m})(x)$ is either of the form
@@
and therefore $|R_{\rho}\mathbf{f}(x)-R_{\rho}\mathbf{f}(x-h)|\lesssim|h|^{\delta}\chi(x)^{\lambda_{0}}$ from which (4.3) now follows. It remains to handle the endpoint case. We claim it suffices to establish that for each sequence $\{\mathbf{f}_{n}\}_{n\in\mathbb{N}}\subset{\mathcal{B}}^{\infty}_{0}$, $R_{\rho}\mathbf{f}_{n}$ has a convergent subsequence in $\mathrm{CMO}(\mathbb{R}^{d})$. Indeed, given such an $\{\mathbf{f}_{n}\}_{n\in\mathbb{N}}$, by a diagonalization argument, we may extract a subsequence $\{\mathbf{f}_{n_{k}}\}_{k\in\mathbb{N}}$ such that for each $\rho_{n}=\frac{1}{n}$, $\{R_{\rho_{n}}\mathbf{f}_{n_{k}}\}_{k\in\mathbb{N}}$ is Cauchy in $\mathrm{CMO}(\mathbb{R}^{d})$. Then, for any $\epsilon>0$ pick $n$ large enough that
$\left\|T-R_{\rho_{n}}\right\|_{\infty,\mathrm{BMO}}<\frac{\epsilon}{3}.$
Such an $n$ exists by the condition $\varepsilon_{z}\to 0$ and (2.8). Then, pick $N$ large enough that for all $i,k\geq N$, $\left\|R_{\rho_{n}}\mathbf{f}_{n_{i}}-R_{\rho_{n}}\mathbf{f}_{n_{k}}\right\|_{\mathrm{BMO}(\mathbb{R}^{d})}<\frac{\epsilon}{3}$. Therefore, by the triangle inequality,
$\left\|T\mathbf{f}_{n_{k}}-T\mathbf{f}_{n_{i}}\right\|_{\mathrm{BMO}(\mathbb{R}^{d})}<\epsilon$
whence $\{T\mathbf{f}_{n_{k}}\}_{k\in\mathbb{N}}$ is Cauchy and has a limit in $\mathrm{CMO}(\mathbb{R}^{d})$. Now it remains to show each $R_{\rho}$ is compact. Applying (4.5) with all $p_{j}=\infty$ implies the same pointwise estimates as above, which implies $R_{\rho}:{\mathcal{B}}^{\infty}_{0}\to{\mathrm{C}}_{\mathrm{v}}(\mathbb{R}^{d})$ and that the family $\{R_{\rho}\mathbf{f}:\mathbf{f}\in{\mathcal{B}}^{\infty}_{0}\}$ is equicontinuous. Therefore, by the Arzela-Ascoli theorem and a diagonalization argument, given a sequence $\{\mathbf{f}_{n}\}_{n\in\mathbb{N}}\in{\mathcal{B}}^{\infty}_{0}$, we can obtain a subsequence such that $\{R_{\rho}\mathbf{f}_{n_{k}}\}_{k\in\mathbb{N}}$ is Cauchy in $\left\|\cdot\right\|_{L^{\infty}([-n,n]^{d})}$ for each $n\in\mathbb{N}$. However, the pointwise estimate for $R_{\rho}\mathbf{f}$ shows that given $\epsilon>0$ we can find $n$ large enough that $\left|R_{\rho}\mathbf{f}_{n_{k}}(x)\right|<\frac{\epsilon}{3}$ for $x$ outside $[-n,n]^{d}$. Combining these two facts with the triangle inequality shows that $\{R_{\rho}\mathbf{f}_{n_{k}}\}_{k\in\mathbb{N}}$ is Cauchy in $\left\|\cdot\right\|_{L^{\infty}}$ which is a stronger norm than $\left\|\cdot\right\|_{\mathrm{BMO}(\mathbb{R}^{d})}$. Finally, recalling that $R_{\rho}\mathbf{f}_{n_{k}}\in{\mathrm{C}}_{\mathrm{v}}(\mathbb{R}^{d})$ establishes that the limit belongs to $\mathrm{CMO}(\mathbb{R}^{d})$. ∎

### v2

@@ A CZ form $\Lambda$ is said to be a compact CZ form if for some (all) $M\geq 1$
By the symmetry in Definitions 3.1 and 3.2 above, if $\Lambda$ is a CZ form or compact CZ form, then so is $\Lambda^{\sigma}$ for any $\sigma\in\mathrm{S}_{m+1}$, defined by (1.5). Therefore, if $T$ is a CZO, so is $T^{*,j}$ for any $j=1,\ldots,m$. It is not too difficult to see that if $\mathsf{W}_{\Lambda}^{M}\to 0$ for some $M$ then the same holds for each $M$.
Remark 3.4.
Compact wavelet forms, as defined in Definition 2.6 are compact CZ forms. Almost orthogonality of elements of the wavelet classes, e.g. [DWW]*Lemma 2.3, and the fact that $\varepsilon_{z}\to 0$ readily shows $\mathsf{W}_{\Lambda}(z)\to 0$ for such forms. Furthermore, since wavelet forms possess cancellation in two inputs, all the symbols $\mathfrak{b}_{j}$ are zero. Compact paraproduct forms, the subject of §5 below, are also compact CZ forms. For these, $\mathsf{W}_{\Lambda}(z)\to 0$ for the same reason as the wavelet forms. Furthermore, $\mathfrak{b}_{0}$ is the paraproduct symbol while $\mathfrak{b}_{j}=0$ for $j=1,\ldots,m$. Our representation theorem (Theorem 2 part C. below) states that in fact all compact CZ forms are a linear combination of forms from these two classes.
@@
Let $T$ be an $m$-linear CZO. We say $T$ is a compact CZO if for each $(\vec{p},r)\in Q$, there exists an extension of $T$, $\bar{T}$, such that $\bar{T}(\mathcal{B}^{\vec{p}})$ is precompact in $L^{r}(\mathbb{R}^{d})$, and at the upper endpoint, there exists an extension $\bar{\bar{T}}$ such that $\bar{\bar{T}}(\mathcal{B}^{\infty})$ is precompact in $\mathrm{CMO}(\mathbb{R}^{d})$. Furthermore, if the weaker condition, for all $(p,r)\in Q_{\mathrm{ref}}$ $T(\mathcal{B}^{{\vec{p}}})$ is precompact in $L^{r}(\mathbb{R}^{d})$, holds, then we say $T$ is a compact CZO in the reflexive range.
@@
An important concept closely related to compact operators is that of a weakly convergent sequence. We give a straightforward sufficient condition for generating such families in $L^{r}(\mathbb{R}^{d})$ parameterized by $z\in Z^{d}$.
Lemma 3.7.
Let $1\leq q_{-}<r<q_{+}\leq\infty$ and $\{f_{z}\in L^{q_{-}}(\mathbb{R}^{d})\cap L^{q_{+}}(\mathbb{R}^{d}):z\in Z^{d}\}$ satisfy
  • (i)
    $\left\|f_{z}\right\|_{L^{r}}$ is uniformly bounded,
  • (ii)
    there exists $\beta>0$ such that $\left\|f_{z}\right\|_{L^{q_{-}}}\lesssim t^{\beta}$, $\left\|f_{z}\right\|_{L^{q_{+}}}\lesssim t^{-\beta}$,
  • (iii)
    for $|x-w|\gtrsim t$, $\left|f_{z}(x)\right|\lesssim t^{-\frac{d}{r}}\chi_{z}(x)$.
Then for any $g\in L^{r^{\prime}}(\mathbb{R}^{d})$, $\lim\limits_{z\to\infty}\left\langle f_{z},g\right\rangle\to 0$.
Proof.
By (i) and a standard density argument, it is enough to prove the conclusion for $g\in{\mathrm{C}}^{\infty}_{0}(\mathbb{R}^{d})$. Suppose ${\mathrm{supp}\,}g\subset B(0,R)$. The main step is the trivial consequence of Hölder’s inequality.
$\left|\left\langle f_{z},g\right\rangle\right|\leq\min\left\{\left\|f_{z}\right\|_{L^{\infty}(B(0,R))}\left\|g\right\|_{L^{1}},\left\|f_{z}\right\|_{L^{q_{-}}}\left\|g\right\|_{L^{q_{-}^{\prime}}},\left\|f_{z}\right\|_{L^{q_{+}}}\left\|g\right\|_{L^{q_{+}^{\prime}}}\right\}.$
(iii) ensures that the first quantity is smaller when $z=(w,t)$, $t\sim 1$, and $|w|$ is much larger than $R$. (ii) controls the latter two quantities when $t$ is small or large. ∎
@@
Let $T$ be an $m$-linear CZO. First, if $T$ is compact in the reflexive range, then for any $\varphi^{1},\ldots,\varphi^{m}\in L^{\infty}$ with compact support, and any $j=0,1,\ldots,m$,
(3.6) $\lim_{z\to\infty}t^{md-\frac{d}{2}}\left\|T^{*,j}(\varphi^{1}_{z},\ldots,\varphi^{m}_{z})\right\|_{L^{2}(\mathbb{R}^{d})}=0,\quad\varphi_{z}^{i}={\mathsf{Sy}}_{z}\varphi^{i}.$
Second, if (3.6) holds, then each $\mathfrak{b}_{j}\in\mathrm{CMO}(\mathbb{R}^{d})$ for $j=0,\ldots,m$, and for each $M\geq 1$,
@@
To prove the first conclusion, set $b_{z}^{j}=t^{md-\frac{d}{2}}T^{*,j}(\varphi^{1}_{z},\ldots,\varphi^{m}_{z})$. Since $T^{*,j}$ is bounded from ×$\times_{j=1}^{m}L^{qm}(\mathbb{R}^{d})\to L^{q}(\mathbb{R}^{d})$ for any $1<q<\infty$ (being a CZO) we have
(3.8) $\left\|b_{z}^{j}\right\|_{L^{q}}\lesssim t^{\frac{d}{q}-\frac{d}{2}}.$
Furthermore,
(3.9) $\left\|b_{z}^{j}\right\|_{L^{2}}^{2}=t^{md-\frac{d}{2}}\left\langle b_{z}^{j},T^{*,j}(\varphi^{1}_{z},\ldots,\varphi^{m}_{z})\right\rangle=\left\langle f^{0}_{z},T(f^{1}_{z},\ldots,f^{m}_{z})\right\rangle,$
for $f^{i}_{z}$ which are appropriate permutations and normalizations of $\varphi^{i}_{z}$ and $b^{j}_{z}$. To assign them precisely, we distinguish between $j=0$ and $j=1,\ldots,m$. In the first case, we can simply take
$f_{z}^{i}=\left\{\begin{array}[]{ll}t^{d(1-\frac{1}{2m})}\varphi_{z}^{i}&i=1,\ldots,m\\ b_{z}^{0}&i=0;\end{array}\right.\quad p_{i}=2m,\ i=1,\ldots,m;\quad r=2.$
And for $j=1,\ldots,m$ then
$f_{z}^{i}=\left\{\begin{array}[]{ll}t^{d(1-\frac{1}{4(m-1)})}\varphi_{z}^{i}&i\not\in\{0,j\}\\ t^{d(1-\frac{1}{4})}\varphi_{z}^{j}&i=0\\ b_{z}^{j}&i=j;\end{array}\right.\quad p_{i}=\left\{\begin{array}[]{ll}4(m-1)&i\not\in\{0,j\}\\ 2&i=j;\end{array}\right.\quad r=4$
In either case, as defined, $\{(f^{1}_{z},\ldots,f^{m}_{z}),z\in Z^{d}\}$ is a bounded set in $L^{\vec{p}}$. Crucially, by Lemma 3.7, $f^{0}_{z}$ converges weakly to zero in $L^{r^{\prime}}(\mathbb{R}^{d})$. In the case $f^{0}_{z}=t^{d(1-\frac{1}{r})}\varphi^{j}_{z}$ the conditions are readily verified for $q_{-}=1$ and $q_{+}=\infty$. The more challenging case, when $f^{0}_{z}=b_{z}^{0}$ (the case $j=0$ and $r=2$), also satisfies the assumptions of Lemma 3.7. Indeed, (i) and (ii) are consequences of (3.8) and (iii) can be verified using the kernel estimate (3.1). Now suppose (3.6) fails. Then, by (3.9), there exists $\varepsilon>0$ and $z_{n}\to\infty$ such that
(3.10) $\left|\left\langle f^{0}_{z_{n}},T(f^{1}_{z_{n}},\ldots,f^{m}_{z_{n}})\right\rangle\right|\geq\varepsilon.$
As $T$ was assumed to be compact from $L^{\vec{p}}\to L^{r}$, there exists a subsequence $\zeta_{n}\to\infty$ such that $\{T(f^{1}_{\zeta_{n}},\ldots,f^{m}_{\zeta_{n}})\}$ converges in $L^{r}$. Combined with the weak convergence to zero of $f^{0}_{\zeta_{n}}$, this contradicts (3.10).
To prove the second statement, Cauchy-Schwarz immediately tells us that (3.7) follows from (3.6). The proof of $\mathfrak{b}_{j}\in\mathrm{CMO}(\mathbb{R}^{d})$ is more delicate. To do so, we define the function $\mathfrak{c}_{j}\in\mathrm{BMO}(\mathbb{R}^{d})$ by $\mathfrak{c}_{j}=T^{*,j}[1,1,\ldots,1]$, where $T[\cdot]$ denotes the now classical extension of $T$ to elements of ×$\times_{j=1}^{m}L^{\infty}(\mathbb{R}^{d})$,
(3.11) $T[f_{1},\ldots,f_{m}]:=\lim_{t\to 0}T(f_{1}\theta_{t},\ldots,f_{m}\theta_{t})-c_{t}$
for suitable constants $c_{t}$ (depending $(m+1)$-linearly on $T,f_{1},\ldots,f_{m}$) with $\lim_{t\to 0}c_{t}=c_{0}\in\mathbb{C}$; see e.g. [GT02]*Lemma 1. Notice that when $f_{j}\in L^{\infty}(\mathbb{R}^{d})\cap\mathcal{S}(\mathbb{R}^{d})$, the definition (3.11) differs from $T(f_{1},\ldots,f_{m})$ but only by a constant. Therefore, when we view (3.11) as an element of $\mathrm{BMO}(\mathbb{R}^{d})$, which consists of equivalence classes modulo constants, there is no contradiction. Clearly $\mathfrak{c}_{j}=\mathfrak{b}_{j}$ in our distributional sense (3.3) as well. It is therefore enough to show $\mathfrak{c}_{j}\in\mathrm{CMO}(\mathbb{R}^{d})$, which is well-known [Uch78] to be equivalent to showing for each $\varepsilon>0$ there exists $M>0$ such that
@@
$\sup_{z\not\in\mathbb{B}_{M}}\mathsf{osc}_{z}(\mathfrak{c}_{j})<\varepsilon,\quad\mathsf{osc}_{z}(f)=\frac{1}{t^{d}}\int_{B(w,t)}\left|f-f_{z}\right|,\quad f_{z}=\frac{1}{t^{d}}\int_{B(w,t)}f.$
Let $\varphi\in L^{\infty}$ satisfy $1_{\{|x|\leq 1\}}\leq\varphi\leq 1_{\{|x|\leq 2\}}$ and, in the sense of (3.11), for any $\zeta\in Z^{d}$, we split $\mathfrak{c}_{j}=E+T[\varphi_{\zeta},\ldots,\varphi_{\zeta}]$, where
$E=T[1-\varphi_{\zeta},1,\ldots,1]+T[\varphi_{\zeta},1-\varphi_{\zeta},1,\ldots,1]+\cdots+T[\varphi_{\zeta},\ldots,\varphi_{\zeta},1-\varphi_{\zeta}].$
Using the kernel estimate (3.2), one can see that for $z=(w,t)$ and $\zeta=(w,at)$, $\mathsf{osc}_{z}(E)\lesssim a^{-\delta}$ for $a\geq 2$. Second, by Hölder’s inequality and the fact that $T$ and $T[\cdot]$ only differ by a constant,
$\mathsf{osc}_{z}(T[\varphi_{\zeta},\ldots,\varphi_{\zeta}])=\mathsf{osc}_{z}(b^{j}_{\zeta})\lesssim_{a}\left\|b^{j}_{\zeta}\right\|_{L^{2}}.$
Therefore, given $\varepsilon>0$, first pick $a$ large enough that $\mathsf{osc}_{z}(E)<\frac{\varepsilon}{2}$. Then, by (3.6), we can pick $M$ large enough that for $\zeta\not\in\mathbb{B}_{M}$, $\mathsf{osc}_{z}(T[\varphi_{\zeta},\ldots,\varphi_{\zeta}])<\frac{\varepsilon}{2}$. But there exists $N\gtrsim_{a}M$ such that if $z=(w,t)\not\in\mathbb{B}_{N}$ then $\zeta=(w,at)\not\in\mathbb{B}_{M}$. ∎
@@
$\left|\Lambda\left(\vec{\psi}_{\mathbf{z}},\phi_{\zeta}\right)\right|\lesssim\left(t^{-md}\right)^{1-\theta}\left[\mathbf{z},\zeta\right]_{\delta_{2}}^{\theta}\min\{1,\mathsf{W}_{\Lambda}^{n}(\zeta)\}^{1-\theta},\quad\mathbf{z}\in\mathbb{B}^{m}_{n}(\zeta),\ \mathbf{z}\geq\zeta.$
We choose $\theta$ close enough to $1$ that $md\theta+\delta_{2}\theta-md=\delta_{1}$. With this specific choice of $\theta$, one can easily verify that $\left(t^{-md}\right)^{1-\theta}[\mathbf{z},\zeta]_{\delta_{2}}^{\theta}=[\mathbf{z},\zeta]_{\delta_{1}}$. Setting $\varepsilon_{\zeta}^{1}=\min\{1,\mathsf{W}^{1}_{\Lambda}(\zeta)\}^{1-\theta}$ handles the case $n=1$. For $n\geq 2$, we factor, with $\rho=\delta_{1}-\eta$,
@@
Therefore, one only needs to verify that for $\mathbf{z}\in\mathbb{B}^{m}_{n}(\zeta)\setminus\mathbb{B}^{m}_{{n-1}}(\zeta)$, the first factor is comparable to $2^{-\rho n}$. The proof is concluded by setting $\varepsilon^{n}_{\zeta}=\min\{1,\mathsf{W}^{n}_{\Lambda}(\zeta)\}^{1-\theta}$. ∎
@@
We recycle the proof of the representation theorem in [DGW] relying on Lemma 2.4 and Lemma 2.5. Let $f_{j}\in{\mathcal{S}}(\mathbb{R}^{d})$, $j=0,1,\ldots,m$ and expand each $f_{j}$ using (2.4) to obtain
@@ 4. Compact wavelet forms
Let $T$ be such that $U(f_{0},\ldots,f_{m})=\left\langle f_{0},T(f_{1},\ldots,f_{m})\right\rangle$ is a compact wavelet form. Then $T$ is a compact CZO.
@@
$T(f_{1},\ldots,f_{m})(x)$ is either of the form
@@
and therefore $|R_{\rho}\mathbf{f}(x)-R_{\rho}\mathbf{f}(x-h)|\lesssim|h|^{\delta}\chi(x)^{\lambda_{0}}$ from which (4.3) now follows. It remains to handle the endpoint case. First note that $R_{\rho}$ is clearly bounded on $\mathcal{B}^{\infty}$. We claim it suffices to establish that for each sequence $\{\mathbf{f}_{n}\}_{n\in\mathbb{N}}\subset{\mathcal{B}}^{\infty}$, $R_{\rho}\mathbf{f}_{n}$ has a convergent subsequence in $\mathrm{CMO}(\mathbb{R}^{d})$. Indeed, given such an $\{\mathbf{f}_{n}\}_{n\in\mathbb{N}}$, by a diagonalization argument, we may extract a subsequence $\{\mathbf{f}_{n_{k}}\}_{k\in\mathbb{N}}$ such that for each $\rho_{n}=\frac{1}{n}$, $\{R_{\rho_{n}}\mathbf{f}_{n_{k}}\}_{k\in\mathbb{N}}$ is Cauchy in $\mathrm{CMO}(\mathbb{R}^{d})$. Then, for any $\epsilon>0$ pick $n$ large enough that
(4.6) $\left\|{(T-R_{\rho_{n}})[\cdot]}\right\|_{\infty,\mathrm{BMO}}<\frac{\epsilon}{3},$
recalling the definition (3.11). Such an $n$ exists by the condition $\varepsilon_{z}\to 0$ and (2.10). Then, pick $N$ large enough that for all $i,k\geq N$, $\left\|R_{\rho_{n}}\mathbf{f}_{n_{i}}-R_{\rho_{n}}\mathbf{f}_{n_{k}}\right\|_{\mathrm{BMO}(\mathbb{R}^{d})}<\frac{\epsilon}{3}$. Therefore, by the triangle inequality, (4.6), and recalling that $R_{\rho}[\mathbf{f_{n}}]$ and $R_{\rho}(\mathbf{f}_{n})$ only differ by a constant,
$\left\|T[\mathbf{f}_{n_{k}}]-T[\mathbf{f}_{n_{i}}]\right\|_{\mathrm{BMO}(\mathbb{R}^{d})}<\epsilon$
whence $\{T[\mathbf{f}_{n_{k}}]\}_{k\in\mathbb{N}}$ is Cauchy and has a limit in $\mathrm{CMO}(\mathbb{R}^{d})$. Now it remains to show each $R_{\rho}$ is compact. Applying (4.5) with all $p_{j}=\infty$ implies the same pointwise estimates as above, which implies $R_{\rho}:{\mathcal{B}}^{\infty}\to{\mathrm{C}}_{\mathrm{v}}(\mathbb{R}^{d})$ and that the family $\{R_{\rho}\mathbf{f}:\mathbf{f}\in{\mathcal{B}}^{\infty}\}$ is equicontinuous. Therefore, by the Arzela-Ascoli theorem and a diagonalization argument, given a sequence $\{\mathbf{f}_{n}\}_{n\in\mathbb{N}}\in{\mathcal{B}}^{\infty}$, we can obtain a subsequence such that $\{R_{\rho}\mathbf{f}_{n_{k}}\}_{k\in\mathbb{N}}$ is Cauchy in $\left\|\cdot\right\|_{L^{\infty}([-n,n]^{d})}$ for each $n\in\mathbb{N}$. However, the pointwise estimate for $R_{\rho}\mathbf{f}$ shows that given $\epsilon>0$ we can find $n$ large enough that $\left|R_{\rho}\mathbf{f}_{n_{k}}(x)\right|<\frac{\epsilon}{3}$ for $x$ outside $[-n,n]^{d}$. Combining these two facts with the triangle inequality shows that $\{R_{\rho}\mathbf{f}_{n_{k}}\}_{k\in\mathbb{N}}$ is Cauchy in $\left\|\cdot\right\|_{L^{\infty}}$ which is a stronger norm than $\left\|\cdot\right\|_{\mathrm{BMO}(\mathbb{R}^{d})}$. Finally, recalling that $R_{\rho}\mathbf{f}_{n_{k}}\in{\mathrm{C}}_{\mathrm{v}}(\mathbb{R}^{d})$ establishes that the limit belongs to $\mathrm{CMO}(\mathbb{R}^{d})$. ∎